OpenStudy (anonymous):

HELP FAST A rectangular sign had a perimeter of 44 feet and a length of 14 feet. A designer shortened the length so that the sign became a square with the same width as the old sign. What was the area of the new sign? A. 256 ft2 B. 64 ft2 C. 112 ft2 D. 196 ft2

OpenStudy (anonymous):

@dan815

OpenStudy (doc.brown):

Old sign 2(14) + 2(width) = 44ft

OpenStudy (doc.brown):

I'd divide it all by 2, but that's just me.

OpenStudy (anonymous):

?

OpenStudy (doc.brown):

\[2(14)+2(w)=44\\\frac{2(14)}{2}+\frac{2(w)}{2}=\frac{44}{2}\\14+w=22\]

OpenStudy (anonymous):

8

OpenStudy (doc.brown):

width is 8, yes

OpenStudy (anonymous):

b

OpenStudy (doc.brown):

Bravo